OpenStudy (anonymous):

What is the slope of the line represented by the equation 4x + 8y = 4

OpenStudy (anonymous):

Rearrange into slope-intercept form(y = mx + b) 4x + 8y = 4 First, subtract 4x to both sides.

OpenStudy (anonymous):

what will it look like in that form?

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

We're solving for 'y'

OpenStudy (anonymous):

ok but with the number what willu it look right

OpenStudy (anonymous):

Subtract 4x to both sides: 8y = -4x + 4 Now divide 8 to both sides, what's -4 / 8?

OpenStudy (anonymous):

4

OpenStudy (holly00d1248):

no divide not subtract

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^

OpenStudy (holly00d1248):

@ribhu

OpenStudy (holly00d1248):

did you get the answer @caseyfaye1998

OpenStudy (holly00d1248):

I think she left @ribhu